OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). Please explain how you got the answer.

OpenStudy (anonymous):

First you get the slope (1+1)/(4-5) or (-1-1)/(5-4) which is -2

OpenStudy (anonymous):

Okay, now what do I do? (BTW, thanks for the help. Power went down due to ice storms so couldn't learn this stuff sooner.)

OpenStudy (anonymous):

Then use point slope form which is y-y1=m(x-x1) so y-1=-2(x-4) and no problem glad to help

OpenStudy (anonymous):

from there use the distributive property which states a(b-c)=a(b)+a(-c) so you do -2(x-4)=-2(x)+ -2(-4) which gives you -2x+8 so now you have the equation y-1=-2x+8

OpenStudy (anonymous):

Okay, can you continue? I'd like to figure this out. Can you just write a giant block so I can read it?

OpenStudy (anonymous):

well after that last step I just did all you have to do is add one to both sides of the equation so the whole equation is this y-1=-2(x-4) y-1=-2x+8 y=-2x+9

OpenStudy (anonymous):

y=-2x+9 is your answer

OpenStudy (anonymous):

thanks for the medal